A cyclist cycles down a hill. At the top of the hill of height 50m he reaches a speed of 50ms-1. Calculate his speed at the bottom of the hill.

1
Expert's answer
2022-03-24T11:00:47-0400

The speed of a cyclist at the bottom of the hill

"v_b=\sqrt{v_t^2+2gh}\\
=\sqrt{50^2+2*9.8*50}=59\:\rm m/s"